# 22footnotetext: To whom correspondence should be addressed

### v6

@@ Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-rank tensors. After decades of research, bonvicini2024irreducible recently achieves an explicit ICT decomposition for $n=5$ with factorial time/space complexity. In this work we, for the first time, obtains decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contractions with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. To the best of our knowledge, this is also the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
@@ 1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approach to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$. Spherical tensors serve as the starting point for EGNNs based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-rank representations can be easily computed using tensor products (specifically, outer products in EGNNs) and contractions, making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
@@
where $\hat{T}^{(n;l;q)}$ is an ICT, ${\rm vec}$ is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $l$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $l$ is sometimes greater than $1$). The decomposition matrices $H^{(n;l;q)}$ should be constructed such that $\hat{T}^{(n;l;q)}$ are irreducible. A formal definition is introduced in Section LABEL:sec:ictdecomp. The challenge lies in the fact that $H^{(n;l;q)}$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently has bonvicini2024irreducible obtained the decomposition matrices for $n=5$. Due to these limitations, state-of-the-art EGNNs cannot freely manipulate high-rank ICTs or can only manipulate those whose weights equal their ranks (zaverkin2024higher).
It has been shown that $l=8$ irreducible representations can greatly benefit the model, but under the spherical basis (liao2023equiformerv2). Therefore, the matrices for high-rank orthogonal ICT decomposition are of great importance. Meanwhile, another crucial question is which linear operations applied to the Cartesian tensors are equivariant, as these operations constitute the design space of EGNNs, in other words, what is the vector space of linear mappings $f$ such that
@@
where $g\in G$, $T\in\left(\mathbbm{R}^{3}\right)^{\otimes n}$, and $\rho:G\to GL\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$ is the matrix representation. This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$. More generally, for different input and output spaces, this equivariant space can be denoted as ${\rm Hom}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes l},\left(\mathbbm{R}^{3}\right)^{\otimes k}\right)$. pearce2023brauer leverages brauer1937algebras’s theorem to find the spanning set of the equivariant space. However, as the rank increases, the number of elements in the spanning set becomes greater than the dimension of the space. Thus, it is also essential to identify a basis to reduce the number of parameters required.
@@
In this work, one of our contributions is the proposal of a non-numerical method for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$ in an affordable setting. Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors. bonvicini2024irreducible proposes a novel method to automate the procedure but does not reduce the complexity. The Reduced Row Echelon Form (RREF) algorithm is also required to run on an $n!\times n!$ matrix. Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis. In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverse change-of-basis can bring us back to Cartesian space. The challenge lies in finding such a change-of-basis matrix. We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices. The change-of-basis matrix converts Cartesian tensors to the direct sum of spherical tensors.
If we treat tensors as vectors, we find that different spherical vectors are separated as distinct elements of the full-length vector. Thus, we can multiply it by a diagonal matrix with only zeros and ones on the diagonal. This operation eliminates the other spherical vectors and retains only the desired ones. The inverted change of basis can then convert the spherical direct product space back to the Cartesian tensor product space. The product of these three matrices—the change-of-basis matrix, the diagonal matrix, and the inverse change-of-basis—forms the decomposition matrix, which decomposes the Cartesian tensor into the desired ICTs. This approach takes less than one second on 28-cores Intel® Xeon® Gold 6330 CPU @ 2.00GHz to compute cases for $n=6$, which is already much faster than the current decomposition for $n=5$ by bonvicini2024irreducible. We also show that this process can be further accelerated. Specifically, we prove that path matrices have orthonormal columns up to multiplicative factors, which we leverage to demonstrate that decomposition matrices can be generated via the product of the path matrices and their transpose, with no inverse operation involved.
@@
Another challenge this work addresses is the construction of bases for the equivariant spaces. With the support of Schur’s lemma, we prove that the bases are formed by matrices that: 1) decompose a vector into its ICT, and 2) map an ICT to another ICT in an isomorphic space (i.e., with the same weight). These matrices are easy to construct in the spherical direct sum space, as their elements lie on the diagonal. Then, our path matrices can change the basis from the spherical space to the Cartesian tensor product space. We further extend our conclusions to linear equivariant operations between different direct sum and tensor product spaces. We demonstrate the capability to construct bases for such general equivariant spaces, including spherical tensor product spaces with different parities. This enables the efficient construction of equivariant layers between different spaces, thereby expanding the design space of EGNNs.
@@
  • We propose a non-numerical method to efficiently obtain high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
@@
  • To the best of our knowledge, this is the first analytic, rather than numerical, method for theoretically obtaining arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases.
@@
$\mathcal{V}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{V}$ $N^{(n;j)}$ The multiplicity of the ICT of rank $n$ and weight $j$
@@
$\hat{T}^{(n;j;q;p)}$ An ICT of rank $n$, weight $j$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{V}}$ The set of all possible paths to decompose a tensor in space $\mathcal{V}$
$H^{(n;j;q)}$ An ICT decomposition matrix to obtain ICTs of rank $n$, weight $j$, and index $q$ $\hat{H}^{(\mathcal{V};j;q)}$ A general ICT decomposition matrix of space $\mathcal{V}$, weight $j$ and index $q$
@@
$P^{(path)}$ A path matrix generated by $path$, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by $path$
@@
$\odot$ Contraction ##\## Number of elements in a set
The following contents are organized as follows. In Section 1, we recall some basic ideas of representation theory, spherical harmonics, tensors, tensor product spaces, and we also introduce the Clebsch-Gordan coefficients and the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-rank ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-rank ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The complexity comparison with bonvicini2024irreducible is in Appendix LABEL:sec:complexityanalysis. The visualizations of ICT decomposition matrices and equivariant bases are shown in Appendix LABEL:sec:visualization. A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure LABEL:fig:mindmap.
@@ 2 Preliminaries
In this section, we recap neccessary mathematical preliminaries.
2.1 Irreducible Representations and Spherical Spaces
We first recall some basic facts. A thorough introduction to group theory can be found in inui2012group and brocker2013representations. We first focus on the $O(3)$ group (3-D rotation and reflection), but we also extend our conclusions to $O(n)$, $SO(n)$, and $SU(n)$ in Section LABEL:sec:discussion. Consider a vector $v\in\mathcal{V}$. The rotation and reflection operations should be able to act on it. Representation theory shows that $O(3)$ can be represented by some real matrices in the Cartesian space via a group homomorphism $\rho:G\to\mathbbm{R}^{3\times 3}$, which is called the natural representation. It holds that
@@ 2.1 Irreducible Representations and Spherical Spaces
for any $g_{1}$, $g_{2}\in G$, and * is the group multiplication. The group element $g$ acting on the vector $v$ can therefore be represented as a matrix-vector multiplication $\rho(g)\cdot v$. Sometimes, one may also refer to the space of $v$ as a representation space. If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{V}$, this space is also called a $G$-invariant space or subspace. To represent $O(3)$, the matrix $\rho(g)$ is constrained to be orthogonal. It is important to note that the co-domain of $\rho$ can differ, meaning that it can also be represented by matrices of other dimensions. In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$. Now, it suffices to define equivariance. We claim that a linear function $f:\mathcal{V}\to\mathcal{V}^{\prime}$ is equivariant if and only if
@@
where we also call $\mathcal{V}^{\prime}$ a $G$-invariant subspace. Conversely, if there is no such subspace $\mathcal{V}^{\prime}$ and the corresponding representation $\rho_{\mathcal{V}^{\prime}}$ on it for a given space $\mathcal{V}$, then we claim that $\rho_{\mathcal{V}}$ is irreducible.

### v7

@@ Abstract
Irreducible Cartesian tensors (ICTs) play a crucial role in the design of equivariant graph neural networks, as well as in theoretical chemistry and chemical physics. Meanwhile, the design space of available linear operations on tensors that preserve symmetry presents a significant challenge. The ICT decomposition and a basis of this equivariant space are difficult to obtain for high-rank tensors. After decades of research, bonvicini2024irreducible has recently achieved an explicit ICT decomposition for $n=5$ with factorial time/space complexity. In this work we, for the first time, obtain decomposition matrices for ICTs up to rank $n=9$ with reduced and affordable complexity, by constructing what we call path matrices. The path matrices are obtained via performing chain-like contractions with Clebsch-Gordan matrices following the parentage scheme. We prove and leverage that the concatenation of path matrices is an orthonormal change-of-basis matrix between the Cartesian tensor product space and the spherical direct sum spaces. Furthermore, we identify a complete orthogonal basis for the equivariant space, rather than a spanning set (pearce2023brauer), through this path matrices technique. Our method avoids the RREF algorithm and maintains a fully analytical derivation of each ICT decomposition matrix, thereby significantly improving the algorithm’s speed to obtain arbitrary rank orthogonal ICT decomposition matrices and orthogonal equivariant bases. We further extend our result to the arbitrary tensor product and direct sum spaces, enabling free design between different spaces while keeping symmetry. The Python code is available at https://github.com/ShihaoShao-GH/ICT-decomposition-and-equivariant-bases, where the $n=6,\dots,9$ ICT decomposition matrices are obtained in 1s, 3s, 11s, and 4m32s on 28-core Intel® Xeon® Gold 6330 CPU @ 2.00GHz, respectively.
@@ 1 Introduction
The deployment of equivariant neural networks in physics, chemistry, and robotics often requires certain equivariance or invariance to effectively incorporate inductive bias. One typical equivariance requirement is equivariance under the $O(3)$ group (rotations and reflections). Equivariant Graph Neural Networks (EGNNs) are well-suited for this purpose, as they can maintain equivariance while processing features within the graph (zaverkin2024higher; simeon2024tensornet; batzner20223). One of the most promising approaches to constructing equivariant layers involves leveraging irreducible representations. Depending on the basis of the space, two distinct types of irreducible representations are utilized in different EGNNs: spherical (spherical tensors) (musaelian2023learning; batatia2022mace) and Cartesian irreducible representations (Irreducible Cartesian Tensors, ICTs) (zaverkin2024higher; schutt2021equivariant) for $O(3)$. Spherical tensors serve as the starting point for EGNNs based on irreducible representations. However, they require Clebsch-Gordan transforms with significant computational overhead to derive desired irreducible representations from existing ones during each inference, making them computationally intensive. In contrast, ICTs are more efficient, as high-rank representations can be easily computed using tensor products (specifically, outer products in EGNNs) and contractions, making them more computationally friendly for most deep learning frameworks. The ICTs are obtained through the ICT decomposition matrices $H$,
@@
where $\hat{T}^{(n;l;q)}$ is an ICT, $\rm vec$} is the vectorization operation via flattening the tensor, $T^{(n)}$ is a given tensor, $n$ represents the rank of the tensor, and $l$ is called the weight (the order of irreducible representations, which is distinct from the learnable weights in neural networks) of the irreducible representation, and $q$ denotes the label (since the multiplicity for given $n$ and $l$ is sometimes greater than $1$). In this paper, we adopt the convention that $\operatorname{vec}$ flattens a tensor by stacking its entries from the rightmost to the leftmost index; concretely, ${\rm vec}(T)=\{T_{...111},T_{...112},\cdots\}$. The decomposition matrices $H^{(n;l;q)}$ should be constructed such that $\hat{T}^{(n;l;q)}$ are irreducible. A formal definition is introduced in Section LABEL:sec:ictdecomp. The challenge lies in the fact that $H^{(n;l;q)}$ of rank $n>5$ are unknown. The recent development of EGNNs research benefits from the irreducible decomposition matrices $H$ obtained from the physics and chemistry communities. The decomposition matrices for $n=4$ have been established by andrews1982irreducible. However, only very recently has bonvicini2024irreducible obtained the decomposition matrices for $n=5$. Due to the lack of high-rank ICT decomposition matrices, it is impossible to construct high-rank Cartesian EGNNs that incorporate valid equivariant linear layers. That is because only irreducible representations of same weight can be linearly combined, but the irreducible representations under Cartesian basis are mixed together. In other words, the irreducible representations of the same weight are impossible to be separated, so we cannot linearly combine those of the same type, which limits the available rank of irreducible representations in EGNNs with full access to the equivariant linear layer designs. The fact that utilizing high-rank irreducible representations improves performance has been well investigated by predecessors, with the support of several empirical evidence. For example, see Table 2 in the paper by batzner20223, where rank $l=3$ NequIP significantly outperforms the same model with lower ranks. On the other hand, in the work done by liao2023equiformerv2, Table 1(c) shows that the performance is better with higher ranks. Therefore, the matrices for high-rank orthogonal ICT decomposition are of great importance. Based on this, we can obtain valid equivariant linear operations applied to the Cartesian tensors, as these operations constitute the design space of EGNNs. In other words, what is the vector space of linear mappings $f$ such that
@@
where $g\in G$, an element of a group $G$. $T\in\left(\mathbbm{R}^{3}\right)^{\otimes n}$ is an arbitrary rank-$n$ tensor. $\rho:G\to GL\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$ is the group representation, and $GL(V)$ is a space consisting of the invertible matrices that operate on vector space $V$. A detailed introduction to the basic and necessary group theory is in Section 2.1. This vector space is called an equivariant space, also denoted as ${\rm End}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes n}\right)$. More generally, for different input and output spaces, this equivariant space can be denoted as ${\rm Hom}_{G}\left(\left(\mathbbm{R}^{3}\right)^{\otimes l},\left(\mathbbm{R}^{3}\right)^{\otimes k}\right)$. pearce2023brauer leverages brauer1937algebras’s theorem to find the spanning set of the equivariant space. However, as the rank increases, the number of elements in the spanning set becomes greater than the dimension of the space. Thus, it is also essential to identify a basis to reduce the number of parameters required.
@@
In this work, one of our contributions is the proposal of a method that does not rely on RREF algorithm, for efficiently obtaining ICT decomposition matrices, through which we achieve $6\leq n\leq 9$ in an affordable setting. Several previous methods (andrews1982irreducible; dincckal2013orthonormal; bonvicini2024irreducible) rely on evaluating the full permutation of $n$ (e.g., fundamental isotropic Cartesian tensors) and $2n$ (e.g., natural projection tensors) indices, which requires factorial complexity, making it impractical for application to high-rank tensors. bonvicini2024irreducible proposes a novel method to automate the procedure but does not reduce the complexity. The Reduced Row Echelon Form (RREF) algorithm is also required to run on an $n!\times n!$ matrix. Instead, we leverage the fact that the Cartesian tensor product space is equivalent to a direct product of spherical spaces, up to a change of basis. In spherical space, it is straightforward to identify the desired irreducible representations, as they are diagonalized; then, the inverse change-of-basis can bring us back to Cartesian space. The challenge lies in finding such a change-of-basis matrix. We prove that one can construct this matrix by concatenating the results of sequential contractions and the outer products of Clebsch-Gordan matrices (CG matrices) according to the parentage scheme (coope1965irreducible), which we refer to as path matrices. The orthogonality of the path matrices further allows us to eliminate the time-consuming inverse operation, rendering this approach an orthogonal decomposition. Both theoretical and experimental analyses confirm the efficiency of this approach. Inspired by Schur’s lemma, which states that only irreducible representations of the same weight can be linearly combined, we construct an equivariant basis via a path-matrix technique. This construction makes it possible to design high-rank equivariant linear layers in EGNNs. We then address the lack of a suitable parentage scheme for general spaces. To this end, we propose a generalized parentage scheme that yields an equivariant basis for arbitrary spaces. These advances collectively enable the efficient construction of equivariant layers across different spaces, thereby expanding the design space of EGNNs.
@@
  • We propose a method to efficiently obtain high-rank ICT decomposition matrices and, for the first time, obtain the decomposition matrices for $6\leq n\leq 9$, significantly improving upon the current limit of $n=5$, as shown in Table 1.
@@
  • Our method avoids the RREF algorithm and derives each ICT decomposition matrix in a fully analytical manner, thereby yielding a significantly more efficient algorithm.
@@
$\mathcal{V}^{\otimes n}$ A rank $n$ tensor product space from $\mathcal{V}$ $N^{(n;l)}$ The multiplicity of the ICT of rank $n$ and weight $l$
@@
$\hat{T}^{(n;l;q;p)}$ An ICT of rank $n$, weight $l$, index $q$, and parity $p$ (sometimes omitted) $Q^{(l;p)}_{\mathcal{V}}$ The set of all possible paths to decompose a tensor in space $\mathcal{V}$
[…]
@@
$P^{(path)}$ A path matrix generated by pathpath, e.g., $(1\to 3\to 2)$ $\hat{P}^{(path)}$ A normalized path matrix generated by pathpath
@@
$\odot$ Contraction #\# Number of elements in a set
The following contents are organized as follows. In Section 2, we recall basic concepts from representation theory, spherical harmonics, tensors, and tensor product spaces, and we introduce the CG coefficients along with the parentage scheme. In Section LABEL:sec:problem, we formulate our problems: obtaining high-rank ICT decomposition matrices and bases for arbitrary tensor product spaces. In Section LABEL:sec:highorderict, we introduce our theory and method to obtain high-rank ICT decomposition matrices. In Section LABEL:sec:basis, we present the theory and method to obtain bases of equivariant spaces where the input and output spaces are the same. In Section LABEL:sec:extension, we extend the theory to the equivariant design spaces where the input and output spaces are different, and show the method to obtain bases for such general spaces. In Section LABEL:sec:relatedworks, we review related works on designing EGNNs and calculating ICT decomposition matrices. In Section LABEL:sec:discussion, we discuss the relation between pearce2023brauer and this work, as well as the theory for $SO(n)$, $O(n)$, and $SU(n)$. We provide concluding remarks in Section LABEL:sec:conclusion, and introduce the reproducibility details in Section LABEL:sec:repro. The acknowledgments are presented after the main text. The proofs of supportive lemmas are shown in Appendix LABEL:sec:proofs. The complexity comparison with bonvicini2024irreducible is in Appendix LABEL:sec:complexityanalysis. We give an example for generating rank-2 ICT decomposition matrices in Appendix LABEL:sec:examplerank2. The visualizations of ICT decomposition matrices and equivariant bases are shown in Appendix LABEL:sec:visualization. A list of mathematical symbols is provided in Table LABEL:tab:symbols, and a mind map of this work is shown in Figure LABEL:fig:mindmap.
@@ 2 Preliminaries
In this section, we recap necessary mathematical preliminaries.
2.1 Group, Irreducible Representations and Spherical Spaces
We first recall some basic facts of group theory. A thorough introduction to group theory can be found in inui2012group and brocker2013representations. A group $G$ is a set equipped with a binary operation *, such that the following five properties hold.
  • Non-emptiness:
    The set $G$ is non-empty.
  • Closure:
    For all $g_{1},g_{2}\in G$, we have $g_{1}*g_{2}\in G$.
  • Associativity:
    For all $g_{1},g_{2},g_{3}\in G$, we have $(g_{1}*g_{2})*g_{3}=g_{1}*(g_{2}*g_{3})$.
  • Identity element:
    There exists $e\in G$ such that, for every $g\in G$, it holds that eg=ge=ge*g=g*e=g.
  • Inverse element:
    For every $g\in G$ there exists $g^{-1}\in G$ satisfying $g*g^{-1}=g^{-1}*g=e$.
Group is a powerful tool to study symmetry. Rotations and rotations with reflections are both examples of groups. For example, it is easy to verify that 2-d rotations satisfy the above requirements: clockwise rotating by $30^{\circ}$, followed by counterclockwise $15^{\circ}$, is equal to clockwise rotating by $15^{\circ}$, and the remained properties can be examined similarly. Until now, the group remains abstract, and we cannot bring it into computation. Representation theory provides us with a bridge to connect group to linear algebra. A representation is a mapping $\rho:G\to GL(V)$ such that
@@ 2.1 Group, Irreducible Representations and Spherical Spaces
holds for any $g_{1},g_{2}\in G$, where $GL(V)$ is a space consisting of the invertible matrices that operate on vector space $V$. In other words, they are invertible matrices that operate on vector space $V$. Here, we introduce groups that we will consider in this work. The orthogonal group, $O(n)$, consisting of orthogonal matrices in $\mathbbm{R}^{n\times n}$, represents $n$-dimensional rotations and reflections; The special orthogonal group, $SO(n)$, consisting of orthogonal matrices with determinant equaling $1$, represents $n$-dimensional rotations; The unitary group, $U(n)$, consisting of unitary complex matrices in $\mathbbm{C}^{n\times n}$, represents $n$-dimensional rotations and reflections on $\mathbbm{C}^{n\times n}$, and $SU(n)$ with additional unit determinant constraint, represents rotations only. Consider a vector $v\in\mathcal{V}$. A group element $g$ acting on the vector $v$ can therefore be represented as a matrix-vector multiplication $\rho(g)\cdot v$. Sometimes, one may also refer to the space of $v$ as a representation space. If the results of the group action remain in the same space, i.e., $\rho(g)v\in\mathcal{V}$, this space is also called a $G$-invariant space or subspace. It is important to note that the co-domain of $\rho$ can differ, meaning that it can also be represented by matrices of other dimensions. In this paper, we denote the co-domain of $\rho$ using a subscript when it should be made explicit, e.g., $\rho_{\mathbbm{R}^{3}}$. Now, it suffices to define equivariance. We claim that a linear function $f:\mathcal{V}\to\mathcal{V}^{\prime}$ is equivariant if and only if
@@
where we also call $\mathcal{V}^{\prime}$ a $G$-invariant subspace. Conversely, if there is no such subspace $\mathcal{V}^{\prime}$ and the corresponding representation $\rho_{\mathcal{V}^{\prime}}$ on it for a given space $\mathcal{V}$ satisfying the above conditions, then we claim that $\rho_{\mathcal{V}}$ is irreducible.